OpenStudy (anonymous):

f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2. 2, 3, 5, 7, 9 2, 3, 4, 7, 11 2, 3, 5, 9, 11 2, 3, 5, 8, 13 @sidsiddhartha

OpenStudy (sidsiddhartha):

and u need to find what?

OpenStudy (anonymous):

Generate the first 5 terms of this sequence:

OpenStudy (sidsiddhartha):

well i think it will produce a fibonacci series

OpenStudy (sidsiddhartha):

put one by one n=3,4,5 and see what u get

OpenStudy (anonymous):

ok

OpenStudy (sidsiddhartha):

so if n=3 then f(n) = f(n - 1) + f(n - 2) f(3)=f(2)+f(1)=2+3=5 like this

OpenStudy (anonymous):

n=3 =5

OpenStudy (sidsiddhartha):

yeah now try n=4

OpenStudy (anonymous):

whats f(3)= to

OpenStudy (sidsiddhartha):

f(3)=5

OpenStudy (anonymous):

ok n=4 =8

OpenStudy (sidsiddhartha):

yeah good now put n=5

OpenStudy (anonymous):

n=5 =13

OpenStudy (sidsiddhartha):

yeah good so what do u think what is the right option?

OpenStudy (anonymous):

last one 2,3,5,8,13

OpenStudy (sidsiddhartha):

correct well done ^_^

OpenStudy (anonymous):

thanks a lot

OpenStudy (sidsiddhartha):

no pro :)

OpenStudy (anonymous):

can yuo help me with another assignment linear models

OpenStudy (sidsiddhartha):

yeah will try

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

A school puts on a play. The play costs $1,200 in expenses. The students charge $4.00 for tickets. There will be one performance of the play in an auditorium that seats 500 people. What is the domain of the function that shows the profit as a function of the number of tickets sold? The domain is all real numbers from negative 1,200 to positive 2,000 The domain is all integers from negative 1,200 to positive 2,000. The domain is the integers from 0 to 500. The domain is all real numbers from 0 to 500.

OpenStudy (anonymous):

sorry i went to eat